# Demonstration of sequential processors with quantum advantage and analysis of classical performance limits

### v1

@@ IV.3 Bound for four-module trit processor
In this section, we considered the problem of how well a classical sequential processor with a single bit or single trit communication can approximate a given target function of distributed input data. The target functions considered by now were specific ones generated by the outputs of quantum processors for the same data. By showing the minimum number of errors in the approximation, we show the gap in expressivity of the classical processor with respect to the quantum processor for specific tasks. The proofs of this section are based on explicit properties of the target functions.
@@
In the next section, we consider a more general problem. Given an arbitrary binary target function, what is the best approximation by an output of a sequential processor? In general, this problem belongs to the high-complexity class. Solving it efficiently would imply an efficient algorithm for, for example, a low-rank binary matrix approximation, which is known to be NP-hard [27, 28, 29]. In the following section, we reduce the problem to an Ising problem, which can be approachable by nowadays and future Ising solvers based on simulated or quantum annealing, or other techniques.
V Optimal classical sequential processors
@@
Consider an $N$ stage sequential processor with one-bit communication between modules and two-bit inputs in each module. The local inputs are not known to other modules. The performance of the processor can be optimized on the strategies of the modules.
@@
    Decide if a given binary target function $T(\{X_{i}\})$ can be achieved as an output from the processor.
@@
    Determine the smallest Hamming distance between the output of the classical processor $O(\{X_{i}\})$ and the target function $T(\{X_{i}\})$.
V.1 Task 1: Existence criterion
For simplicity, assume that the number of processing modules is $N=3$. We can answer the first challenge by checking the number of different blocks of size 4, and 16 in the target function. There are 16 blocks of size 4. They are indexed by sequences of bits $(X_{1}X_{2}X_{3}X_{4})$ and defined as
$t^{[4]}(X_{1}...X_{4},X_{5}X_{6})=T(X_{1},...,X_{6}).$
There are 4 blocks of size 16. They are indexed by sequences of bits $(X_{1}X_{2})$ defined as
$t^{[16]}(X_{1}X_{2},X_{3}...X_{6})=T(X_{1},...,X_{6})$
Theorem 1. The processor defined above can output functions $O(X_{1},...,X_{6})$ with at most two different rows in the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$ and at most two different rows in the reshuffled matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$.
Proof.
There are four possible outputs (for each $X_{1}X_{2}$, one) from the first processor but only two binary symbols 0 or $1$. For each symbol there is one and only one succession (the row of the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$), as the following processor cannot distinguish the same symbols output for different inputs of the first processor. So, there can be only two different rows in matrix $O^{[4]}$. Similarly, there are 16 outputs (for each sequence $(X_{1}...X_{4})$, one) from the second processor, but only two binary symbols. For each symbol, there is one succession. The row of the matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$ for a chosen value of $(X_{1}...X_{4})$ is equal to one of these successions. If we recognize that the target function consists of only two distinct rows $O^{[16]}$ and only two distinct rows $O^{[4]}$ we can associate symbols 0 or $1$ as the output of the first processor in the positions corresponding to different distinct rows. Similarly, we associate 0 or $1$ as the outputs of the second processor in the positions corresponding to distinct rows of $O^{[16]}$. ∎
The argument can be easily extended to processors with more modules and different limits on communication between the modules.
Therefore, we can formulate the exisence criterion as the following theorem.
Theorem 2. [Existence criterion] For a binary function of several variables, one can check if it can be an output from a sequential processor with a single bit communication and a given distribution of variables between the modules by checking if each reshuffling of this function into a matrix with rows indexed by the local variables of earlier modules, and columns indexed by the remaining variables is row-rank at most 2.
@@ V.1 Task 1: Existence criterion
For sequential processors with a single trit communication, the row-rank should be 3, etc.
V.2 Task 2: Optimal circuit as the Ising problem
V.2.1 Example
@@
Figure 6: Two stage sequential processor with one bit communication and one bit input $X_{i}$ at each stage. The output function of inputs is denoted by $O(X_{1},X_{2})$.
@@
For simplicity, let us consider an example – a two-module sequential processor with one bit communication and one bit input at each stage, as shown in Fig. 6. The first part, depending on the bit $X_{1}$ it obtains as the input, chooses which binary variable to pass to the next stage. Its strategy is given by a binary function $F_{1}$ which is defined by binary variables $a_{1}$ and $a_{2}$ returned for the respective input $X_{1}$. The second module strategy $F_{2}$ is a function of the bit passed from the previous part and one extra bit $X_{2}$. This function is given by a four-bit binary vector of binary variables $a_{3},...,a_{6}$, see Fig. 7. The goal is to choose the values of the variables $a_{1},...,a_{6}$ such that the output from the last part of the processor for given variables $X_{1}$ and $X_{2}$ was as close as possible to the given binary target function $T(X_{1},X_{2})$.
@@
Figure 7: Example of binary variables which define each part’s strategy functions $F_{i}$ of their respective inputs. In this there are two parts in the processor. Each part processes the bit obtained from the previous part and one extra bit $X_{i}$.
@@
For a two-stage processor, the problem is trivial because the target has only 4 values. We can associate the first two elements of the target function with $a_{3}$ and $a_{4}$, and the other two with $a_{5}$ and $a_{6}$ and put $a_{1}=0$ and $a_{2}=1$. However, we want to show that we can obtain this result as the optimization of an Ising-like problem. Then we generalize the entire formalism for any number of stages and more general setups in which the problem is highly nontrivial.
For any $X_{1}$ and $X_{2}$ the output of the processor $O(X_{1},X_{2})$ is
$O(X_{1},X_{2})=(1-a_{1+X_{1}})a_{3+X_{2}}+a_{1+X_{1}}a_{5+X_{2}}$ (12)
@@
For binary functions, the minimum Hamming distance coincides with the maximum correlation function, which in this example is given as
@@
$+$ $T(1,1)[(1-a_{2})a_{4}+a_{2}a_{6}]$
The last formula is a quadratic form in binary variables, which can be written as an Ising problem with non-local two-body interactions. Its maximum provides the set of bits $a_{1},...,a_{6}$ that maximizes $C$.
V.2.2 Generalization
The construction of the example above can easily be generalized to arbitrary processor length, arbitrary communication, and arbitrary number of local input bits. Let us give here the explicit formula for a three-stage processor with a one-bit communication and a two-bit local input, as in the lower part of Fig. 3. The binary variables defining the strategies are shown in Fig. 8.
Figure 8: Example of binary variables which define each part’s strategy functions $F_{i}$ of their respective inputs. In this example there are three modules in the processor. Each part processes a bit obtained from the previous part and two extra bits.
The formula for the corelations between the target and the output reads
@@ V.2.2 Generalization
In this case, the correlation function is a polynomial of degree three of binary variables which can be mapped in a standard way to the Ising model with up to three spins and possibly far-range interactions.
@@
Further generalizations are needed to consider four-stage processors and the processor which allow for trit-based communication. In this case, variables $a_{i}$ are integer numbers from the set $\{0,1,2\}$. Because routinely we use solvers acting on binary variables, before formulating the formulas for correlations we first map each trit-variable to two one-bit variables. We do not show the final formulas that can be easily obtained with the same strategies as described above.
V.2.3 Solvers
@@
In the previous sections, we show that finding the highest correlations between the output of the classical sequential processor with restricted communication and a given binary target function is equivalent to solving an Ising problem for a chain of spins with nonlocal and multi-spin interactions. This problem is, in general, NP-complete, and efficient solvers for every case do not exist. One uses approximate solvers applying specific heuristics.
@@
In our research, we used a dedicated computer provided by Fixstars Amplify [30]. It finds solutions through the process of simulated annealing. We used the solver to find candidate solutions for the maximum correlation between the output of the classical three-module sequential processor with two-bit local variables and one-bit communication and the target given by the deterministic part of the output of the corresponding quantum processor with one-qubit communication. The problem involved 20 binary variables (spins) as shown in Fig. (8). We defined the Ising Hamiltonian corresponding to the correlation function (V.2.2). The solver was able to find the solution with 8 errors with respect to the target function. It is shown in Table 1. Proposition 1 states that this is the lower bound on the number of errors for this target. Therefore, we conclude that the solver found the optimal solution.
@@
We repeated the numerical experiment on Fixstars Amplify solver successfully finding cases corresponding to the optimal strategies of the classical processors with a single trit communication, three- and four-modules and the target functions given by the deterministic parts of the corresponding quantum optical processors with qutrit communication as in the quantum optics experiment. The target functions are given (11) or the first four rows of it. The lower bounds on the number of errors are given in Propositions 2 and 3. As the optical processor could deterministically achieve only two solutions, $|20\rangle$ and $|02\rangle$, the number of bits in the output of the last stage was the same as in the bit processor, which is 4 for strategies associated with possibly three different outputs from the previous module. The total number of binary variables in the optimization of the three-module trit processor was 44. We found the strategies of each module that achieved the minimum number of errors which is shown in Table 2.
@@
For finding the optimal four-module trit processor strategy with Fixstars Amplify, we needed (for our license restriction) to limit the number of binary variables to 48. We did it by guessing the strategies of the first and last modules. The solver found that the optimal strategies of the intermediate modules as shown in Table 3.
VI Extensions and Outlook

### v2

@@ II.2 Quantum one-qutrit processor
As proven in Appendix A, for the three-stage trit processor the classical bound on the correlation with the target function is 0.4375. This corresponds to at least two errors between the output of the optimized processor and the non zero output of the target function. The explicit strategy of such processor is shown in Table 2. For the four-stage processor, the classical limit on the correlation with the above target is 0.375, which corresponds to at least 16 errors in the optimized output. This strategy is shown in Table 3.
III Silicon photonics experiment
We implemented both a one-qubit and a one-qutrit processor on a silicon photonic integrated circuit [17, 18, 19] and evaluated their performance using the correlation function described in Sec. II. As shown in Fig. 3(a), the experimental setup employed a $4\times 4$ universal unitary transformation circuit composed of beam splitters and phase shifters [20, 21]. In this study, we used two spatial modes of the circuit to construct a sequential processor by cascading multiple unitary transformations. Each unitary transformation consisted of a phase shifter and a Mach–Zehnder interferometer, allowing us to realize the target operations with high precision.
Figure 3: (a) Schematic of the sequential quantum processor implemented on a silicon photonic chip. The unitary gate was implemented using a heater and an Mach–Zehnder interferometer enclosed by the red dotted lines. (b) Overview of the one-qubit processor. A heralded single photon was injected into the sequential processor. (c) Overview of the one-qutrit processor. Two photons generated from attenuated laser light at the single-photon level were injected into the sequential processor. In panels (b) and (c), the circuit enclosed by the red dotted lines corresponds to a single unitary gate in each case.
For the implementation of the one-qubit processor, heralded single photons were employed, as illustrated in Fig. 3(b). Photon pairs were generated via spontaneous four-wave mixing (SFWM), utilizing the third-order nonlinearity of silicon [22, 23]. Specifically, a continuous-wave (CW) pump light with a wavelength of 1549.32 nm and an output power of approximately 10 mW was injected into a silicon waveguide. This generated signal and idler photons through the SFWM. The signal photons had a central wavelength of 1559.0 $\pm$ 0.8 nm, while the idler photons had a central wavelength of 1539.8 $\pm$ 0.8 nm. The spectral width of these photons is determined by the bandwidth of the off-chip frequency filter (200 GHz). The signal photons were used as heralds (triggers), and the idler photons were directed to the processor circuit [24, 25, 26]. After passing through these filters, both the idler and the corresponding signal photons were detected using superconducting nanowire single-photon detectors. The generation rate of signal single photons obtained after passing through the circuit was approximately 30 counts per second.
In this experiment, the input state was $|10\rangle_{ab}$, where a single photon was present in mode $a$. Randomly selected operations from the set $\{\hat{I},\hat{\sigma}_{z},\hat{H},\hat{H}^{*}\}$ were applied through three successive unitary transformations. At the output, the occurrence probabilities of the states $|10\rangle_{ab}$ and $|01\rangle_{ab}$ were measured from the photon number distribution obtained over a 1-s integration period. Using approximately 85000 detection events, the correlation function with respect to the target state was calculated.
Figure 3 shows a histogram of the obtained correlation function. To construct the histogram, the approximately 85000 experimental data points were divided into subsets of 1000 events each. The correlation function was calculated for each subset, and the frequency distribution of these values was plotted. From the histogram, the mean of the correlation function was 0.490, and the square root of the variance (standard error of the mean) was 0.017. These results indicate that the experimentally obtained correlation function significantly exceeds the limit of 0.25 achievable by the classical processor with one-bit communication. In the figure, the vertical blue line corresponds to this limit.
Next, for the implementation of the one-qutrit processor, we attenuated a laser beam to the single-photon level and approximated a two-photon state by post-selecting events in which two photons were detected simultaneously [Fig. 3(c)]. Specifically, each of the two output ports of the photonic integrated circuit was connected to an off-chip fiber beam splitter, followed by single-photon detectors at the two outputs of each splitter. Coincidence detections of one photon at each detector were used to identify events corresponding to the $\lvert 20\rangle_{ab}$ or $\lvert 0;2\rangle_{ab}$ states. This detection configuration functions as a pseudo-photon-number-resolving scheme and allows us to detect photon bunching in each mode.
To suppress multi photon contributions, the input laser was strongly attenuated such that the single-count rate was approximately $2.5\times 10^{5}$ counts per $0.5~\mathrm{s}$. Under this condition, we observed approximately $5000$ coincidence events after the beam splitter (corresponding to about $10000$ events before the splitter). The mean photon number of the attenuated coherent state was estimated to be $\sim 0.08$. Given that the photon-number statistics follow a Poisson distribution, the contribution of three-photon or higher-order events to the measurement results is estimated to be $\sim 0.008\%$. Therefore, the impact of multi photon components introduces a negligibly small error of below $0.01\%$ in the experimental data.
The input state for the qutrit experiment was $|20\rangle_{ab}$, corresponding to two photons in mode $a$. Four unitary transformations, randomly selected from the operations defined in Eqs. (1)–(4) and (10)–(13), were sequentially applied. The output probabilities of the states $|20\rangle_{ab}$ and $|02\rangle_{ab}$ were measured, and approximately 65,000 detection events were collected. Correlation functions were then calculated based on these data.
Figure 4 shows a histogram of the correlation function for a four-module quantum processor with one qutrit communication. As in Fig. 3, approximately 65000 experimental data points were divided into subsets of 1000 events each, and the correlation function was calculated for each subset. The histogram displays the frequency distribution of the resulting correlation values with respect to the target function (14). From the histogram, the mean of the correlation function was 0.525, and the square root of the variance was 0.014. These results indicate that the experimentally obtained correlation function significantly exceeds the classical limit which, as shown in Fig. 5, is 0.375.
Figure 4: The histogram of the correlation function based on the data from three-module single-qubit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation, which for the considered target function is equal to 0.25.
Figure 5: The histogram of the correlation function based on the data from four-module single-qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation, which for the considered target function is equal to 0.375.
Figure 6: The histogram of the correlation function based on the data from three-module single-qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.4375.
Finally, in post-processing, from the data of the qutrit experiment with four modules, we extracted data that correspond to the first gate fixed as the identity. In this way, we performed the analysis of the three-module processor with single-qutrit communication. The target function in this case is given by the first four rows of Eq. (14). We built the histogram of the values of the correlation function for 65 subsets of data with approximately 250 points in each subset. The average is 0.509 with the standard error of the average 0.068. This allows us to conclude that the observed value of the correlation significantly exceeds the limit for the correlation achievable by the classical processor with one-trit communication, which, as shown in Fig. 6, is equal to 0.4375.
In conclusion, in each case that we analyzed, we observe that the realistic experiment produces the correlation function close to 0.5, which is the maximum value of the correlation function with respect to the target [Eqs. (8) or (14)] with half of the entries equal to zero. The bounds on the achievable correlation function for the optimal classical sequential processors are shown in Figs. 46 as the vertical blue lines are established in Section II. We conclude that the classical processor cannot output a function that approaches the target function for any deterministic strategy. In this sense, the quantum processor significantly outperforms the classical one in this task, showing a separation of expressivity between the classical and quantum processors.
IV Optimal classical sequential processors
@@
In the previous sections, we considered the problem of how well a classical sequential processor with single-bit or single-trit communication can approximate a given target function of distributed input data. The target functions considered until now were specific ones generated by the outputs of quantum processors. We showed the gap in expressivity of the classical processor with respect to the quantum processor for specific tasks. The proofs derived in Appendix A are based on explicit properties of the target functions.
In this section, we consider a broader problem. Given an arbitrary binary target function, we ask what is the best approximation that can be obtained by the output of a classical sequential processor. The result can be useful to establish rigorous bounds on classical processors in other designs and experimental demonstrations of quantum advantage in limited space processors.
The problem what is the class of function which can be obtained as an output of a quantum processor depends on specific constraints. We will not answer this question in this paper. However, establishing bounds on classical processor of a given structure is a first step to identify the classes of functions achievable in quantum regime, as we know that all processors that are classical in ideal conditions can be realized by the quantum counterparts.
@@
In general, saying how close a given binary function of Boolean arguments can be approximated by outputs of a classical processor with a fixed structure belongs to a class of high complexity. To solve it efficiently, one would need an algorithm akin to solving the low-rank binary matrix approximation problem [27, 28, 29]. However, this problem or its alternatives are known to be NP hard. In this section, we reduce this task to an Ising problem, which can be approachable by nowadays and future Ising solvers based on simulated or quantum annealing, or other techniques and, under certain conditions, can be efficiently solved.
@@
Consider an $N$-stage sequential processor with one-bit communication between parts and two-bit inputs in each module. The local inputs are not known to other parts. The performance of the processor can be optimized on the strategies of the modules.
@@
    Task 1: Decide if a given binary target function $T(\{X_{i}\})$ can be achieved as an output from the processor.
@@
    Task 2: Determine the smallest Hamming distance between the output of the classical processor $O(\{X_{i}\})$ and the target function $T(\{X_{i}\})$.
IV.1 Task 1: Existence criterion
We formulate the existence criterion as the following practical procedure proven in Appendix B. Consider a binary function $T$ of several variables. Consider also an arbitrary sequential processor with a single-bit communication and a given distribution of variables $\{X_{i}\}$ between the modules. The following procedures allow us to check if the sequential processor can output $T$ on $\{X_{i}\}$:
  1. 1.
    Reshuffle the function $T$ into a matrix $M$ whose rows are indexed by the local variables of earlier modules and columns are indexed by the remaining variables.
  2. 2.
    Compute the row rank of $M$.
  3. 3.
    Repeat for other reshufflings of $T$.
If, for any $M$, the associated row rank is at most 2, $T$ can be outputted by the sequential processor.
@@ IV.1 Task 1: Existence criterion
For sequential processors with single-trit communication, the row rank should be 3, and so on.
IV.2 Task 2: Optimal circuit as the Ising problem
IV.2.1 Example
@@
Figure 7: Two-stage sequential processor with one-bit communication and one bit input $X_{i}$ at each stage. The output function of inputs is denoted by $O(X_{1},X_{2})$.
@@
For simplicity, let us consider an example—a two-module sequential processor with one-bit communication and one-bit input at each stage, as shown in Fig. 7. The first part, depending on the bit $X_{1}$ it obtains as the input, chooses which binary variable to pass to the next stage. Its strategy is given by a binary function $F_{1}$ which is defined by binary variables $a_{1}$ and $a_{2}$ returned for the respective input $X_{1}$, i.e.,
$F_{1}(X_{1})=\begin{cases}a_{1}\text{ if }X_{1}=0,\\ a_{2}\text{ if }X_{1}=1.\end{cases}$
The second module strategy $F_{2}$ is a function of the bit passed from the previous part and an extra bit $X_{2}$. This function’s output is given by a four-bit binary vector of binary variables $a_{3},...,a_{6}$ (see Fig. 8). The goal is to choose the values of the variables $a_{1},...,a_{6}$ such that the output from the last part of the processor, for given variables $X_{1}$ and $X_{2}$, was as close as possible to the given binary target function $T(X_{1},X_{2})$.
@@
Figure 8: Example of binary variables that define each part’s strategy functions $F_{i}$ of their respective inputs. In this case, there are two parts in the processor. Each part processes the bit obtained from the previous part and one extra bit $X_{i}$.
@@
For a two-stage processor, the problem is trivial because the target has only four values. We can associate the first two elements of the target function with $a_{3}$ and $a_{4}$, and the other two with $a_{5}$ and $a_{6}$ and put $a_{1}=0$ and $a_{2}=1$. However, we want to show that we can obtain this result as the optimization of an Ising like problem. Then, we generalize the entire formalism to any number of stages and more general setups in which the problem is highly nontrivial.
For any $X_{1}$ and $X_{2}$, the output of the processor $O(X_{1},X_{2})$ is
$O(X_{1},X_{2})=(1-a_{1+X_{1}})a_{3+X_{2}}+a_{1+X_{1}}a_{5+X_{2}}.$ (15)
@@
For binary functions, the minimum Hamming distance coincides with the maximum of the correlation function. In this example, it is given as
@@
$+$ $T(1,1)[(1-a_{2})a_{4}+a_{2}a_{6}].$
The last formula is a quadratic form in binary variables. This can be written as an Ising problem with nonlocal two-body interactions. Its maximum provides the set of bits $a_{1},...,a_{6}$ that maximizes $C$.
IV.2.2 Generalization
The construction of the example above can easily be generalized to arbitrary processor length, arbitrary communication, and arbitrary number of local input bits. Let us give here the explicit formula for a three-stage processor with one-bit communication and a two-bit local input, as in the lower part of Fig. 4. The binary variables that define the strategies are shown in Fig. 9.
Figure 9: Example of binary variables that define each part’s strategy functions $F_{i}$ of their respective inputs. In this example, there are three modules in the processor. Each part processes a bit obtained from the previous part and two extra bits.
The formula for the correlations between the target and the output reads
@@ IV.2.2 Generalization
In this case, the correlation function is a polynomial of degree three in the binary variables that can be mapped in a standard way to the Ising model with up to three spins, and possibly far-range interactions.
@@
Further generalizations are needed to consider four-stage processors and the processor that allows for trit-based communication. In this case, variables $a_{i}$ are integers from the set $\{0,1,2\}$. Because we routinely use solvers acting on binary variables, before formulating the formulas for correlations, we first map each trit variable to two one-bit variables. We do not show the final formulas that can be easily obtained with the same strategies as described above.
IV.2.3 Solvers
@@
In the previous sections, we showed that finding the highest correlations between the output of the classical sequential processor with restricted communication and a given binary target function is equivalent to solving an Ising problem for a chain of spins with nonlocal and multi-spin interactions. This problem is NP hard, and efficient solvers, for the general case, do not exist. However, one can use specific heuristics based on approximate Ising solvers.
@@
In our research, we use a dedicated computer provided by Fixstars Amplify [30]. It finds solutions through the process of simulated annealing. We used the solver to find candidate solutions for the maximum correlation. The correlation is taken between the output of the classical three-module sequential processor with two-bit local variables and one-bit communication and the target given by the deterministic part of the output of the corresponding quantum processor with one-qubit communication. The problem involved 20 binary variables (spins) as shown in Fig. 9. We defined the Ising Hamiltonian corresponding to the correlation function (IV.2.2). (The number of terms in the Hamiltonian is roughly the number of nonzero elements in the target function times 2 to the power of the number of modules. Note that this is much less than the number of all possible binary functions which for one-bit processor with two-bit inputs in each module is roughly 2 to the power of eight times the number of modules. The brute force optimization would require this number of steps.) The solver was able to find the solution with eight errors with respect to the target function. as shown in Table 1. As proven in Appendix A, this is the lower bound on the number of errors for this target. Therefore, we conclude that the solver found the optimal solution.
@@
We repeated the numerical experiment on Fixstars Amplify solver successfully finding cases corresponding to the optimal strategies of the classical processors with single-trit communication, three and four modules, and the target functions given by the deterministic parts of the corresponding quantum optical processors with qutrit communication as in the quantum optics experiment. The target functions are given in Eq. (14) or the first four rows of it. The lower bounds on the number of errors are derived in Appendix A. As the optical processor could deterministically achieve only two solutions, $|20\rangle$ and $|02\rangle$, the number of bits in the output of the last stage was the same as in the one-bit processor. This is 4 for strategies associated with possibly three different outputs from the previous module. The total number of binary variables in the optimization of the three-module trit processor was 44. We found strategies of each module that achieved the minimum number of errors, which is shown in Table 2.
@@
For finding the optimal four-module trit processor strategy with Fixstars Amplify, we needed (for license restriction) to limit the number of binary variables to 48. We did it by guessing the strategies of the first and last modules. The solver found the optimal strategies of the intermediate modules as shown in Table 3.
V Discussion: extension of the results